A manufacturer ships toasters in cartons of 50. In each​ carton, they estimate a 4​% chance that one of the toasters will need to be sent back for minor repairs. In a batch of 25,000 ​toasters, what are the chances that fewer than 950 need to be​ returned?

The probability that fewer than 950 toasters need to be returned is nothing.
Answer #1

Ans:

mean=0.04*25000=1000

std. dev=sqrt(25000*0.04*0.96)=30.98

z=(950-1000)/30.98

z=-1.61

P(z<-1.61)=0.0534